Derive the formula of Wnetmax in an air standard Brayton Cycle by maximizing the formula of Wnet in terms of T2 and using the relationship T1/T2=T4/T3 . Your answer should be as a function of m, Cp, T1, T2 and T3 only. Using the derived formula, find Wnetmax when T1= 300C, T3=1000C and T2=500C in kJ/kg.
Answer #1

Solution-

Assumption-

  • Kinetic and potential energies across turbine and compressor neglected.
  • Steady state operation
  • Air Standard assumptions valid.

Let us consider a brayton cycle with 1-2 as Isentropic compression .2-3 Heat addition in combustion chamber , 3-4 Isentropic expansion in turbine , 4-1 Heat rejection to atmosphere at constant pressure.

Analysis-

1 Qin Wt. 4 like Qout. L S We = (kl) comp E During process Ito 2 Compression vosk input. inco (T2-T) kJ. Buring process ? toh

2 dlwhet) - mico micg {0+ m2 -1 +0] =- d (T2) Tits = 12 T2 = √ToT3. using Relationship in equation @ Tu Tito o for max Work n

Result-

From the above relationship for (W net)max is a function of T3 = T max temperature which is constraint by material properties of gas turbine . T1 = T min which is constraint by atmosphere(ambient temperature and pressure).

Condition for (W net)max

\small T_{2}=T_{4}= \sqrt{T_{1}*T_{3}}